Parker Plastic, Inc., manufactures plastic mats to use with rolling office chairs. Its standard cost information for last year follows: Standard Quantity Standard Price (Rate) Standard Unit Cost Direct materials (plastic) 12 sq ft. $ 0.72 per sq. ft. $8.64 Direct labor 0.25 hr. $ 12.20 per hr. 3.05 Variable manufacturing overhead (based on direct labor hours) 0.25 hr. $ 1.20 per hr. 0.30 Fixed manufacturing overhead ($378,000 ÷ 900,000 units) 0.42 Parker Plastic had the following actual results for the past year: Number of units produced and sold 1,000,000 Number of square feet of plastic used 11,800,000 Cost of plastic purchased and used $ 8,260,000 Number of labor hours worked 245,000 Direct labor cost $ 2,891,000 Variable overhead cost $ 318,500 Fixed overhead cost $ 355,000 Required: Calculate Parker Plastic’s direct materials price and quantity variances. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable.)
Answer #1

Material price variance = (Standard price-actual price)actual quantity purchased = (0.72*11800000-8260000) = 236000 F

Material quantity variance = (Standard qty-actual qty)Standard price = (1000000*12-11800000)0.72 = 144000 F